# Moments and the Basic Equilibrium Equations

### v1

@@ Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition provides equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so some dynamical equilibrium configurations, such as relative equilibria, may be treated with our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
@@
Affine Geometry, Barycentric Coordinates, Statics, $N$-Body Problem, Central Configurations.
@@ 1. Introduction
Equilibrium problems have contributed to the development of Geometry and Physics since ancient Greece. The work of Archimedes on the principle of the lever is considered the beginning of the mathematical treatment of Statics, but even before him, there had appeared books attributed to Aristotle and to Euclid on the dynamics and balance of the lever, respectively [15]. The concept of barycenter, or center of gravity, of a system of point masses, and its application to studying equilibrium configurations go back to Archimedes. As we shall review in section 2, the barycenter is defined for sets of weighted points whose total weight, $\mu_{0}$, is nonzero as the unique point where the first moment, $\mu_{1}$, vanishes. In the 17th century, C. Huygens used $\mu_{1}$ to introduce the notion of expected value in probability, and realized the importance of the second moment, $\mu_{2}$, in the form of the moment of inertia (Euler’s terminology), for the analysis of pendulum motion. Later on G. W. Leibniz and J. S. König applied $\mu_{2}$ in Geometry and Dynamics [19]. A consequential event was the invention by A. F. Möbius of homogeneous, barycentric coordinates in the early 19th century [13]. In contemporary Mathematical Statistics and Analysis, among other areas, one finds the classical (Stieltjes, Hausdorff, Hamburger) moment problem: given a sequence of moments of a real random variable, determine the (unique) probability distribution (measure) that produced the sequence [27, 30].
@@
Central configurations is a research topic with a long history and recognized relevance to Celestial Mechanics [4, 3, 28, 31, 34]. One of the reasons is that the only known explicit solutions of the Newtonian $n$-body problem, namely the homographic motions, start at such configurations (cf. Wintner [34], §375). The possible homographic motions include the relative equilibria, which are planar periodic solutions, and, in recent times, have been shown to occur in any Euclidean space of even dimension [3, 28]. In the context of the $n$-body problem, it is common to use the terminologies planar (or two-dimensional) central configurations and relative equilibria interchangeably. Notwithstanding the many particular questions that have been answered, so far a classification of central configurations is known only for three arbitrary masses, thanks to Euler [17] and Lagrange [26], and for four equal masses, thanks to Albouy [2], who based his work on a previously obtained symmetry result by himself [1]. General, basic problems on central configurations remain unsolved. An outstanding example is the celebrated Smale’s sixth problem for the 21st century [32, 33], which asks whether the number of similarity classes of relative equilibria of $n$ positive point masses is finite for all $n$. In the broader sense of central configurations of arbitrary dimensions, Smale’s problem is part of the Chazy-Wintner-Smale finiteness conjecture [4]. The finiteness problem has been fully solved only for $n=4$, by Hampton and Moeckel [22]. Considering just generic finiteness with respect to the masses, finiteness has been proved for $n=5$ in dimension two by Albouy and Kaloshin [5], and for the central configurations of $n$ bodies in dimension $n-2$, the so-called Dziobek configurations, by Moeckel [29] (see also Dias [14]). Below we consider the most important sets of central configuration equations used in the classification and finiteness problems, interspersed with summaries of our present contributions.
@@
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek where expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension $n-2$, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
@@
Notably the whole Dziobek formalism is contained in the theory of weighted systems $(X,w)$ with $\mu_{1}=\overrightarrow{O}$. The weight functions of such systems (with fixed $X$) forms a vector space, $\mathbb{W}_{0}(X)$, and we describe a method for constructing bases for $\mathbb{W}_{0}(X)$ using volume barycentric coordinates. The case $\dim\mathbb{W}_{0}(X)=1$ of our theory, and the fundamental hypothesis (FC) of page 1, are implicit in Albouy’s algebraic approach: in essence, it is the ideas in reference [4] that we recast and generalize to $\dim\mathbb{W}_{0}(X)>1$. Besides the intensive use of barycentric coordinates, this generalization requires some additional concepts, in particular that of codimension of a configuration, which plays a chief role throughout the remainder of this paper. A significant number of equilibrium (including central configuration) equations is introduced through the “codimension-exterior power formula”
@@
where $c=n-d-1$ is the codimension of $X$, $n$ is the number of points in $X$, and $d$ is dimension of the affine closure of $X$, usually referred to as the dimension of $X$. It would not be an overstatement to say that (1.1) synthesizes the basic theory of systems with $\mu_{1}=\overrightarrow{O}$. Formula (1.1) is used in section 5 to provide a broad generalization of some determinantal identities obtained by Dziobek [16, 21] for central configurations with codimension one. Besides, (1.1) yields tailored equilibrium equations that exploit properties such as the symmetry and the dimension of a configuration or its subconfigurations, as illustrated in example 5.3. In view of its versatility, we consider identity (1.1) one of the main contributions of our work. We also remark that the Dziobek-type central configuration equations of Dias [14] are here extended to arbitrary weighted systems with $\mu_{1}=\overrightarrow{O}$, and we show that they actually provide a characterization of such systems having any specific dimension. Finally, a somewhat surprising connection with Algebraic Geometry arises as we interpret barycentric coordinates as Plücker coordinates: Plücker relations [24, 25] for Grassmannians of $d$-dimensional subspaces of $\mathbb{R}^{n-1}$ emerge directly from the generalization of Dias’ formulæ. We remark that Plücker relations, the linear relations for volume barycentric coordinates which form our basis vectors of $\mathbb{W}_{0}(X)$, and the generalized Dziobek equations of section 6 may be used to formulate equilibrium problems entirely in terms of barycentric coordinates, via corollary 2.6.
@@
Regarding applications, it is also worth noting that all the weight functions defined by the models of interacting particles having $X$ as an equilibrium configuration coinhabit the space $\mathbb{W}_{0}(X)$. Thus they are a fortiori interrelated, and their relationships may be expressed through equations such as those produced by formula (1.1). In this context, our bases for $\mathbb{W}_{0}(X)$ supply a purely geometric (i.e. non-physical) common framework for representing the weight functions of each individual model.
@@
Going back to the topic of central configurations, we remark that Dziobek equations were also used in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems actually employed the set of central configuration equations introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration, which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
@@
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the article we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the fundamental hypothesis (FC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension, or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure any given dimension is expressed in terms of the configuration’s codimension $c$ simply as (${c+1\choose 2}$. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], §$\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
@@
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in our work, namely, assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, for each $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine each $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.7). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action $=$ $(-1)$$\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and (${c+1\choose 2}$ parameters, where $c$ is the codimension of $X$.
@@
The fundamentals of the theory of moments of weighted systems are reviewed in section 2, culminating with the Huygens-Leibniz-König (HLK) theorem, which collects classical results relating $\mu_{0},\mu_{1}$ and $\mu_{2}$. We demonstrate three corollaries of the HLK theorem with conditions on the second moment $\mu_{2}$ of a weighted system with $\mu_{0}=0$ to have $\mu_{1}=\overrightarrow{O}$. Prototypes of the generalized Albouy-Chenciner equations and the extended Leibinz identities appear at this point. Together with barycentric coordinates, which are introduced in section 4, the HLK theorem and its corollaries are the main tools for our analysis of equilibrium problems.
Equilibrium equations for systems of interacting particles are firstly discussed in section 3. According to our fundamental hypothesis (FC), the total force on each particle $x\in X$ is a linear combinations with coefficients $\varphi_{x,y}$ of the relative positions from $x$ to the remaining particles $y\in X\setminus\{x\}$. We apply the results of section 2 to obtain equilibrium equations that are either linear or bilinear on the $\varphi_{x,y}$, and linear on the squared mutual distances between points of $X$. These equations hold independently of the dimension of $X$ and contain, in particular, the generalization of the Albouy-Chenciner equations and the extended Leibniz identities.
Section 4 is dedicated to the set of all weighted systems with a fixed configuration $X$ and $\mu_{1}=\overrightarrow{O}$. The corresponding weight functions form a vector space $\mathbb{W}_{0}(X)$ whose dimension is the codimension of $X$. Equation (1.1) immediately follows. In section 4, we view Dziobek configurations as simplexes $S$ with an additional point in the affine closure of $S$. This geometric interpretation, together with the associated volume barycentric coordinates, allow us to introduce devices called Dziobek trees in order to construct explicit bases for $\mathbb{W}_{0}(X)$. Section 4 closes with a set of homogeneous second degree equations that characterize the weight functions in $\mathbb{W}_{0}(X)$. Remark 4.10 discusses applications to Grassmannians.
Section 5 concerns applications of the previous sections to central configurations. After providing a formal definition of central configurations followed by a short review of basic results, some classical formulæ are reobtained, and new central configuration equations are discussed. We consider the not so explored problem of central configurations with zero total mass, for which some geometric properties and algebraic relations are directly determined. In addition, we examine the planar five-body central configurations having three collinear bodies, assuming the total mass is nonzero. Our results impose some significant restrictions on the possible central configurations which are consistent with previous studies of particular cases by other authors. Appendix B describes yet another set of central configuration formulæ.
@@
The article concludes with two appendices. In appendix A we show how the HLK theorem of section 2 can be used to find constraints for mutual distances. We apply corollary 2.6 to obtain constraints corresponding to either fixing the dimension of the configuration $X$, which take the form of vanishing Cayley-Menger determinants of Dziobek subconfigurations of $X$, or imposing co-sphericity, which are expressed as the vanishing of the determinant of the so-called relative configuration matrix [3, 28] of $X$. Two theorems are proved in appendix A. In theorem A.2.1 we clarify the relationship between the space of weight functions with $\mu_{1}=\overrightarrow{O}$, $\mathbb{W}_{0}(X)$, and the kernels of the relative configuration matrix and the Cayley-Menger matrix of $X$. The dimension of the configuration is thus related to vanishing Cayley-Menger determinants of subconfigurations. Then, in theorem A.2.2, we obtain the expression (${c+1\choose 2}$ for the number of independent constraints that ensures a configuration has a fixed dimension. Finally, in appendix B, we determine the kernel of the relative configuration matrix viewed, as in reference [3], as a bilinear form. The extended Leibniz identities follow therefrom.

### v2

@@ Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition provides equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
@@
Affine Geometry, Barycentric Coordinates, Equilibria, $N$-Body Problem, Central Configurations.
@@ 1. Introduction
Equilibrium problems have contributed to the development of Mathematics and Physics since ancient Greece. The work of Archimedes on the principle of the lever is considered the beginning of the mathematical treatment of Statics, but even before him, there had appeared books attributed to Aristotle and to Euclid on the dynamics and balance of the lever, respectively [15]. The concept of barycenter, or center of gravity, of a system of point masses, and its application to studying equilibrium configurations go back to Archimedes. As we shall review in section 2, the barycenter is defined for sets of weighted points whose total weight, $\mu_{0}$, is nonzero as the unique point where the first moment, $\mu_{1}$, vanishes. In the 17th century, C. Huygens used $\mu_{1}$ to introduce the notion of expected value in probability, and realized the importance of the second moment, $\mu_{2}$, in the form of the moment of inertia (Euler’s terminology), for the analysis of pendulum motion. Later on G. W. Leibniz and J. S. König applied $\mu_{2}$ in Geometry and Dynamics [19]. A consequential event was the invention by A. F. Möbius of homogeneous, barycentric coordinates in the early 19th century [13]. In contemporary Mathematical Statistics and Analysis, among other areas, one finds the classical (Stieltjes, Hausdorff, Hamburger) moment problem: given a sequence of moments of a real random variable, determine the (unique) probability distribution (measure) that produced the sequence [27, 30].
@@
Central configurations is a research topic with a long history and recognized relevance to Celestial Mechanics [4, 3, 28, 31, 34]. One of the reasons is that the only known explicit solutions of the Newtonian $n$-body problem, namely the homographic motions, start at such configurations (cf. Wintner [34], §375). The possible homographic motions include the relative equilibria, which are planar periodic solutions, and, in recent times, have been shown to occur in any Euclidean space of even dimension [3, 28]. In the context of the $n$-body problem, it is common to use the terminologies planar (or two-dimensional) central configurations and relative equilibria interchangeably. Notwithstanding the many particular questions that have been answered, so far a classification of central configurations is known only for three arbitrary masses, thanks to Euler [17] and Lagrange [26], and for four equal masses, thanks to Albouy [2], who based his work on a previously obtained symmetry result by himself [1]. General, basic problems on central configurations remain unsolved. An outstanding example is the celebrated Smale’s sixth problem for the 21st century [32, 33], which asks whether the number of similarity classes of relative equilibria of $n$ positive point masses is finite for all $n$. In the broader sense of central configurations of arbitrary dimension, Smale’s problem is part of the Chazy-Wintner-Smale finiteness conjecture [4]. The finiteness problem has been fully solved only for $n=4$, by Hampton and Moeckel [22]. Considering just generic finiteness with respect to the masses, finiteness has been proved for $n=5$ in dimension two by Albouy and Kaloshin [5], and for the central configurations of $n$ bodies in dimension n2n-2, the so-called Dziobek configurations, by Moeckel [29] (see also Dias [14]). Below we consider the most important sets of central configuration equations used in the classification and finiteness problems, interspersed with summaries of our present contributions.
@@
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek where expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension n2n-2, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
@@
Notably the whole Dziobek formalism is contained in the theory of weighted systems $(X,w)$ with $\mu_{1}=\overrightarrow{O}$. The weight functions of such systems (with fixed $X$) form a vector space, $\mathbb{W}_{0}(X)$, and we describe a method for constructing bases for $\mathbb{W}_{0}(X)$ using volume barycentric coordinates. The case $\dim\mathbb{W}_{0}(X)=1$ of our theory, and the fundamental hypothesis (FC) of page 1, are implicit in Albouy’s algebraic approach: in essence, it is the ideas in reference [4] that we recast and generalize to $\dim\mathbb{W}_{0}(X)\geq 1$. Besides the intensive use of barycentric coordinates, this generalization requires some additional concepts, in particular that of codimension of a configuration, which plays a chief role throughout the remainder of this paper. A significant number of equilibrium (including central configuration) equations is introduced through the “codimension-exterior power formula”
@@
where c=nd1c=n-d-1 is the codimension of $X$, $n$ is the number of points in $X$, and $d$ is dimension of the affine closure of $X$, usually referred to as the dimension of $X$. It would not be an overstatement to say that (1.1) synthesizes the basic theory of systems with $\mu_{1}=\overrightarrow{O}$. Formula (1.1) is used in section 5 to provide a broad generalization of some determinantal identities obtained by Dziobek [16, 21] for central configurations with codimension one. These identities yield tailored equilibrium equations that exploit properties such as symmetry and dimension, as illustrated in example 5.3. In view of its versatility, we deem equation (1.1) one of the main contributions of our work. We also remark that the Dziobek-type homogeneous, quadratic central configuration equations of Dias [14] are here extended to arbitrary weighted systems with $\mu_{1}=\overrightarrow{O}$, and we show that they actually provide a characterization of such systems having any specific dimension. Finally, a somewhat surprising connection with Algebraic Geometry arises as we interpret barycentric coordinates as Plücker coordinates: Plücker relations [24, 25] for Grassmannians of $d$-dimensional subspaces of $\mathbb{R}^{n-1}$ emerge directly from the generalization of Dias’ formulæ.
@@
Regarding applications, two remarks seem in order. Firstly, Plücker relations, the linear relations for volume barycentric coordinates which form our basis vectors of $\mathbb{W}_{0}(X)$, and the generalized Dziobek equations of section 6 may be used to formulate equilibrium problems entirely in terms of barycentric coordinates, via corollary 2.6. And secondly, all the weight functions defined by the models of interacting particles having $X$ as an equilibrium configuration coinhabit the space $\mathbb{W}_{0}(X)$. Thus they are a fortiori interrelated, and their relationships may be expressed through equations such as those produced by formula (1.1). In this context, our bases for $\mathbb{W}_{0}(X)$ supply a purely geometric (i.e. non-physical) common framework for representing the weight functions of each model individually.
@@
Going back to the topic of central configurations, it is worth noting that Dziobek equations also found an application in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems actually employed a different set of central configuration equations, namely the one introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration, which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
@@
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the paper we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the weaker fundamental hypothesis (WFC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension, or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure the configuration has a given dimension is expressed in terms of the configuration’s codimension $c$ simply as $c+1\choose 2$}. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], $\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
@@
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, for each $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine the $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.7). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action == (1)(-1) $\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and $c+1\choose 2$} parameters, where $c$ is the codimension of $X$.
@@
The fundamentals of the theory of moments of weighted systems are reviewed in section 2, culminating with the Huygens-Leibniz-König (HLK) theorem, which collects classical results relating $\mu_{0},\mu_{1}$ and $\mu_{2}$. We demonstrate three corollaries of the HLK theorem with conditions on the second moment $\mu_{2}$ of a weighted system with $\mu_{0}=0$ to have $\mu_{1}=\overrightarrow{O}$. Prototypes of the generalized Albouy-Chenciner equations and the extended Leibinz identities appear at this point. Together with barycentric coordinates, which are introduced in section 4, the HLK theorem and its corollaries are the main tools for our formulations of equilibrium problems.
Equilibrium equations for systems of interacting particles are firstly discussed in section 3. According to our fundamental hypothesis (FC), the total force on each particle $x\in X$ is a linear combinations with coefficients $\varphi_{x,y}$ of the relative positions from $x$ to the remaining particles $y\in X\setminus\{x\}$. We apply the results of section 2 to obtain equilibrium equations that are either linear or bilinear on the $\varphi_{x,y}$, and linear on the squared mutual distances between points of $X$. These equations contain, in particular, the generalization of the Albouy-Chenciner equations and the extended Leibniz identities.
Section 4 is dedicated to the set of all weighted systems with a fixed configuration $X$ and $\mu_{1}=\overrightarrow{O}$. The corresponding weight functions form a vector space $\mathbb{W}_{0}(X)$ whose dimension is the codimension of $X$. Equation (1.1) immediately follows. In section 4, we view Dziobek configurations as simplexes $S$ with an additional point in the affine closure of $S$. This geometric interpretation allow us to introduce devices called Dziobek trees for $X$ in order to construct explicit bases for $\mathbb{W}_{0}(X)$. Section 4 closes with a set of homogeneous second degree equations that characterize the weight functions in $\mathbb{W}_{0}(X)$. Remark 4.10 discusses applications to Grassmannians.
Section 5 concerns some illustrative applications of the previous sections to central configurations. After providing a formal definition of central configurations followed by a short review of basic results, some classical formulæ are reobtained, and new central configuration equations are discussed. We consider the not so explored problem of central configurations with zero total mass, for which some geometric properties and algebraic relations are directly determined. In addition, we examine the planar five-body central configurations having three collinear bodies, assuming the total mass is nonzero. Our results impose some significant restrictions on the possible central configurations which are consistent with previous studies of particular cases by other authors.
@@
The article concludes with two appendices. In appendix A we show how the HLK theorem of section 2 can be used to find constraints for mutual distances. We apply corollary 2.6 to obtain constraints corresponding to either fixing the dimension of the configuration $X$, which take the form of vanishing Cayley-Menger determinants of Dziobek subconfigurations of $X$, or imposing co-sphericity, which are expressed as the vanishing of the determinant of the so-called relative configuration matrix [3, 28] of $X$. Two theorems are proved in appendix A. In theorem A.2.1 we clarify the relationship between the space of weight functions with $\mu_{1}=\overrightarrow{O}$, $\mathbb{W}_{0}(X)$, and the kernels of the relative configuration matrix and the Cayley-Menger matrix of $X$. The dimension of the configuration is thus related to vanishing Cayley-Menger determinants of subconfigurations. Then, in theorem A.2.2, we obtain the expression $c+1\choose 2$} for the number of independent constraints that ensures a configuration has a fixed dimension. Finally, in appendix B, we determine the kernel of the relative configuration matrix viewed, as in reference [3], as a bilinear form. The extended Leibniz identities follow therefrom, and a final set of central configuration formulæis provided.